OpenStudy (anonymous):

Find the number that this series converges on: 9/17+3/17+1/17+1/51+... Write the number as a fraction in lowest terms, if the answer is a fraction. Do NOT convert to decimals. Write fractions like this: 1/2

OpenStudy (zzr0ck3r):

you need to find the ratio so divide the 2nd term by the first

OpenStudy (zzr0ck3r):

@sweetpink365 what is the ratio? \[ratio = \frac{2nd\space term}{1st \space term}\]

OpenStudy (anonymous):

Im not sure

OpenStudy (zzr0ck3r):

what is the first term?

OpenStudy (zzr0ck3r):

what is the 2nd term?

OpenStudy (zzr0ck3r):

what is the 2nd term divided by the first term

OpenStudy (zzr0ck3r):

if you cant answer this, then you need too stop and go back and figure out how to divide numbers

OpenStudy (anonymous):

The first term is .52941

OpenStudy (zzr0ck3r):

\[\frac{\frac{3}{17}}{\frac{9}{17}}=?\]

OpenStudy (zzr0ck3r):

don't write these in decimal.

OpenStudy (zzr0ck3r):

the first term is 9/17

OpenStudy (anonymous):

so its 2/17 divided by 9/17?